# The Structure of the Majorana Clifford Group

### v2

@@ VI Conclusions
In this paper, we have compiled a comprehensive overview of Majorana strings and the corresponding Clifford operators through the lens of their symplectic representation. Assuming parity superselection, we could identify a subgroup of parity-preserving Cliffords, which is represented by binary orthogonal matrices. We also showed how these p-Cliffords can be efficiently generated from a small set of braiding operators, and how they can be used to prepare any even-parity stabilizer code.
A frame potential analysis of the p-Clifford subgroup restricted to a fixed-parity sector of the Hilbert space showed that it is equivalent to the ordinary Clifford group acting on the same subspace. The p-Clifford group is thus also a 3-design.
One application of the methods developed in this paper is the study of fermionic random tensor network models of SYK as envisioned in [21] and perhaps other fermionic tensor network models of quantum gravity [30]. The results presented here might also be useful in analyzing other Majorana-based systems, such as in fermionic quantum codes and fermionic random circuits, and could perhaps be combined with fermionic tensor network machinery [31]. It could also be interesting to investigate fermionic analogs of the Wigner function, which is a useful tool for characterizing qudit stabilizer state [23], although there are known difficulties in the qubit case which may be inherited in the fermionic context.
@@ VII Acknowledgments
Appendix A Proofs
A.1 Theorem 1
@@
    Using the fact that $\operatorname{\gamma}$ is unitary, one can rewrite (13) into the equivalent statement
    $\begin{split}&[\operatorname{\gamma}\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}][\operatorname{\gamma}\operatorname{\mu}(v^{\prime})\operatorname{\gamma}^{\dagger}]\\ ={}&(-1)^{\left\langle v,v^{\prime}\right\rangle}[\operatorname{\gamma}\operatorname{\mu}(v^{\prime})\operatorname{\gamma}^{\dagger}][\operatorname{\gamma}\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}]\\ \Longleftrightarrow\quad\quad&\operatorname{\mu}(S(v))\operatorname{\mu}(S(v^{\prime}))\\ ={}&(-1)^{\left\langle v,v^{\prime}\right\rangle}\operatorname{\mu}(S(v^{\prime}))\operatorname{\mu}(S(v)),\end{split}$ (88)
    where we used the fact that $c(v)\neq 0$. But since (13) must also hold when we directly set $v\rightarrow S(v)$, we get the following requirement:
@@
    $\frac{1}{|\mathbb{F}_{2}^{2n}|}\sum_{v\in\mathbb{F}_{2}^{2n}}|\operatorname{tr}(\operatorname{\mu}(Sv))|^{2}=1.$ (91)
    Both representations are therefore equivalent and hence (up to an overall sign) connected by a unitary transformation $\gamma$ that induces $S$.
@@
    $\begin{split}&[\operatorname{\mu}(a)\operatorname{\gamma}(S)]\operatorname{\mu}(v)[\operatorname{\mu}(a)\operatorname{\gamma}(S)]^{\dagger}\\ ={}&\operatorname{\mu}(a)\operatorname{\gamma}(S)\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}(S)\operatorname{\mu}(a)\\ ={}&c(v)\cdot\operatorname{\mu}(a)\operatorname{\mu}(Sv)\operatorname{\mu}(a)\\ ={}&(-1)^{\left\langle Sv,a\right\rangle}c(v)\cdot\operatorname{\mu}(Sv)\\ \equiv{}&c^{\prime}(v)\cdot\operatorname{\mu}(Sv),\end{split}$ (92)
@@
    $\left[\frac{c(v)}{c^{\prime}(v)}\right]\left[\frac{c(v^{\prime})}{c^{\prime}(v^{\prime})}\right]=\frac{c(v+v^{\prime})}{c^{\prime}(v+v^{\prime})}$ (93)
    for all $v,v^{\prime}\in\mathbb{F}_{2}^{2n}$. This together with $c(v)^{2}=c(0)=1$ (and the symplectic form being non-degenerate) implies that one can write without loss of generality
    $\frac{c(v)}{c^{\prime}(v)}=(-1)^{\left\langle v,a^{\prime}\right\rangle}=(-1)^{\left\langle Sv,a\right\rangle},$ (94)
    where we set $a^{\prime}\equiv S^{-1}a\in\mathbb{F}_{2}^{2n}$. This coincides with the form of the phase induced by an additional Majorana string $\operatorname{\mu}(a)$, as shown in (LABEL:eq:cliff_majo_action).
@@
    $\begin{split}&\operatorname{\gamma}(S)\operatorname{\gamma}(T)\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}(T)\operatorname{\gamma}^{\dagger}(S)\\ \propto{}&\operatorname{\gamma}(S)\operatorname{\mu}(Tv)\operatorname{\gamma}^{\dagger}(S)\\ \propto{}&\operatorname{\mu}(STv)\\ \propto{}&\operatorname{\gamma}(ST)\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}(ST)\quad\forall v\in\mathbb{F}_{2}^{2n}.\end{split}$ (95)
@@ A.2 Theorem 4
We first show that $\lvert\operatorname{tr}(\Pi_{+}U)\rvert^{2}$ with $U=\operatorname{\mu}(a)\operatorname{\gamma}(S)\in\mathcal{G}$ is either zero or equal to $(f_{+}(S)+c_{+}(S))/2$. For that we evaluate the expression as a sum over the Bell basis $\lvert\Phi_{n}(v)\rangle\equiv(\operatorname{\mu}(v)\otimes I)\lvert\Phi_{n}^{+}\rangle$ of $\operatorname{\mathcal{H}}\otimes\operatorname{\mathcal{H}}$, where $\lvert\Phi_{n}^{+}\rangle\equiv\lvert\Phi^{+}\rangle^{\otimes n}$ denotes $n$ copies of canonical Bell state:
@@
$\begin{split}&\lvert\operatorname{tr}(\Pi_{+}U)\rvert^{2}\\ ={}&\operatorname{tr}(\Pi_{+}U\otimes\Pi_{+}U^{\dagger})\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}\langle\Phi_{n}(v)\rvert\Pi_{+}U\otimes\Pi_{+}U^{\dagger}\lvert\Phi_{n}(v)\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\Pi_{+}U\operatorname{\mu}(v)\otimes\Pi_{+}U^{\dagger}\lvert\Phi_{n}^{+}\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\Pi_{+}U\operatorname{\mu}(v)U^{\dagger}\Pi_{+}\otimes I\lvert\Phi_{n}^{+}\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}c(v)\,\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\Pi_{+}\operatorname{\mu}(Sv)\Pi_{+}\otimes I\lvert\Phi_{n}^{+}\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}c(v)\,\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\operatorname{\mu}(Sv)\Pi_{+}\otimes I\lvert\Phi_{n}^{+}\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{c(v)}{2^{n}}\,\operatorname{tr}[\operatorname{\mu}(v)\operatorname{\mu}(Sv)\Pi_{+}]\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{c(v)}{2^{n+1}}\operatorname{tr}[\operatorname{\mu}(v)\operatorname{\mu}(Sv)+\operatorname{\mu}(v)\operatorname{\mu}(Sv)\operatorname{\mu}(j)]\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{c(v)}{2}(\delta_{Sv,v}+\zeta(j,v)\,\delta_{Sv,v^{c}})\\ \end{split}$ (97)
We can now analyze the structure of $c(v)$ for both cases:
@@
  • In the case of $Sv=v$, it follows directly from (46) that the coefficients of any two (even-parity) fixed points $v,v^{\prime}$ satisfy $c(v)\,c(v^{\prime})=c(v+v^{\prime})$. This together with $c(v)^{2}=1$ implies that we can choose $c(v)=(-1)^{a^{T}v}$ without loss of generality and in accordance with (LABEL:eq:cliff_majo_action).
  • In the case of $Sv=v^{c}$, for any two (even-parity) complemented points $v,v^{\prime}$ (46) can we rewritten as
    $c(v)\,c(v^{\prime})=\frac{\zeta(v,v^{\prime})}{\zeta(v^{c},(v^{\prime})^{c})}\cdot c(v+v^{\prime}).$ (98)
@@
    But since the sum of two complemented points is always a fixed point, we can identify $c(v+v^{\prime})=(-1)^{a^{T}(v+v^{\prime})}$, as before. Additionally, by evaluating $\operatorname{\mu}(j)\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})\operatorname{\mu}(j)$ in two different orders one arrives at
    $\zeta(v,v^{\prime})=\zeta(j,v)\,\zeta(j,v^{\prime})\,\zeta(v^{c},(v^{\prime})^{c}).$ (99)
@@
    Plugging this all into (98) and using $\zeta(j,v)^{2}=1$ (since $\left\langle j,v\right\rangle=0$), results in
    $\left[c(v)\,\zeta(j,v)\right]\left[c(v^{\prime})\,\zeta(j,v^{\prime})\right]=(-1)^{a^{T}(v+v^{\prime})},$ (100)
@@
Applying these results to the evaluation of $\lvert\operatorname{tr}(\Pi_{+}U)\rvert^{2}$ then leaves us with

### v3

@@ VI Conclusions
In this paper, we have compiled a comprehensive overview of Majorana strings and the corresponding Clifford operators through the lens of their symplectic representation. Assuming parity superselection, we were able to identify the subgroup of parity-preserving Cliffords, which is represented by binary orthogonal matrices. We also showed how these p-Cliffords can be efficiently generated from a small set of braiding operators (and the parity operator), and how they can be used to prepare any even-parity stabilizer code (assuming access to an ancilla system composed of two Majorana modes).
Using this framework, a frame potential analysis of the p-Clifford subgroup restricted to a fixed-parity sector of the Hilbert space showed that it is equivalent to the ordinary Clifford group acting on the same subspace. The p-Clifford group interpreted as a random ensemble is thus also a 3-design under the imposed restrictions.
As mentioned in the introduction, one future application of the methods developed in this paper is the study of random tensor network models describing the SYK model, which is envisioned in [22]. But we suspect that other fermionic tensor network models or matchgate circuits (with ties to holography) could also benefit from the toolset established here [36, 37, 38].
More generally, our results might prove to be useful for the in-depth analysis of generic classes of systems, such as fermionic quantum codes, (non-)Gaussian states, random circuits and classical shadows [39]. In that context it could also prove useful to better understand how the presence of a Hamiltonian can impose restrictions on the possible (Jordan-Wigner) transformations [26], and what relative phase differences can be induced by p-Cliffords when acting on superpositions of Majorana strings and quantum states.
Last but not least, it would be interesting to investigate potential fermionic analogs of the Wigner function or the stabilizer Rényi entropy, which are useful tools for characterizing the stabilizerness of bosonic quantum states [23, 40]. But at least for the former there are known difficulties in the qubit case which may be inherited by the fermionic setting.
@@ VII Acknowledgments
Appendix A Canonical Jordan-Wigner Transformation
Based on the $\mathbb{F}_{2}^{2n}$ representation $W$ of the Jordan-Wigner transformation (29), we provide a specific canonical choice in terms of the actual Hilbert space operators. This was done by reading off the columns of $W$ and plugging them into definitions (4) and (24) respectively, thereby implicitly fixing the signs of the mapping. For the latter we also employed the fact that $iXZ=Y$. Using the short-hand tensor product notation
$\begin{split}P_{k}&\equiv I^{\otimes(k-1)}\otimes P\otimes I^{\otimes(n-k)}\end{split}$ (91)
for $P=X,Y,Z$ and $k=1,\ldots,n$, the resulting transformations are then as follows:
Majorana to Pauli:
$\begin{split}\chi_{2k-1}&=Y_{1}Y_{2}\cdots Y_{k-1}\,X_{k},\\ \chi_{2k}&=Y_{1}Y_{2}\cdots Y_{k-1}\,Z_{k}.\end{split}$ (92)
Pauli to Majorana:
$\begin{split}X_{k}&=(i)^{(k-1)\text{ mod }2}\cdot\chi_{1}\chi_{2}\cdots\chi_{2k-2}\,\chi_{2k-1},\\ Z_{k}&=(i)^{(k-1)\text{ mod }2}\cdot\chi_{1}\chi_{2}\cdots\chi_{2k-2}\,\chi_{2k}.\end{split}$ (93)
Appendix B Majorana Algebras
Any element $O$ of a Majorana algebra acting on a $2^{n}$-dimensional Hilbert space $\operatorname{\mathcal{H}}$ can be expressed as a linear combination of Majorana strings, such that
$O\equiv\sum_{v\in\mathbb{F}_{2}^{2n}}c_{v}\cdot\operatorname{\mu}(v),\quad c_{v}\in\mathbb{C}.$ (94)
In fact, it is straightforward to show that a Majorana algebra composed of $2n$ modes encompasses all complex-valued $2^{n}\times 2^{n}$ matrices [28]. Using (34) and the fact that Hermitean strings square to the identity, the coefficients $c_{v}$ of the expansion can be determined using:
$c_{v}=\frac{1}{2^{n}}\operatorname{tr}(O\operatorname{\mu}(v)).$ (95)
where $n_{A}=|A|/2$, and $\Pi$ is some fixed permutation matrix. Note however that we run into issues with regard to the Hilbert space representation of the subalgebra. The strings as depicted in (96) are still endowed with the ordinary $2^{n}\times 2^{n}$-representation (34), even though the algebra itself is isomorphic to the $2^{n_{A}}\times 2^{n_{A}}$ algebra of complex matrices. Hence this representation is reducible, which is especially problematic when considering properties like entropy that intrinsically depend on the Hilbert space dimension. To compensate, one has introduce a new irreducible representation $\operatorname{\mu}_{A}(v_{A})$ which satisfies (4) and (34) with $n\rightarrow n_{A}$. The reduced operator $O_{A}$ corresponding to the chosen subalgebra is then given by
Conversely, one can define the algebraic tensor product $\otimes_{f}$ of two algebras linearly in terms of its action on individual strings:
$\operatorname{\mu}_{A}(v_{A})\otimes_{f}\operatorname{\mu}_{B}(v_{B})\equiv\operatorname{\mu}_{AB}(v_{A}\oplus v_{B}),$ (98)
where $\operatorname{\mu}_{A},\operatorname{\mu}_{B}$ and $\operatorname{\mu}_{AB}$ represent the irreducible representations of their respective algebras. Note however that this operation is only really well-defined for even-parity strings, as otherwise the two possible orders in which to take the product will differ by an overall sign.
Appendix C Examples
C.1 Fermionic Permutation Operators
Weight-2 braiding operators (58) act as simple permutation operators. To see this, let for example:
$\Pi_{ij}\coloneqq\exp\left(-\frac{\pi}{4}\chi_{i}\chi_{j}\right)\equiv\operatorname{BRAID}(i\chi_{i}\chi_{j})\quad(i\neq j).$ (99)
Then from (LABEL:eq:braiding_action) (or direct computation) it follows that the non-trivial action of $\Pi_{ij}$ on single Majorana modes is fully characterized by
$\Pi_{ij}\chi_{i}\Pi_{ij}^{\dagger}=\chi_{j},\quad\Pi_{ij}\chi_{j}\Pi_{ij}^{\dagger}=-\chi_{i}.$ (100)
In that sense, $\Pi_{ij}$ is the equivalent of the SWAP operator for qubits, and occurs in the context of e.g. computing Rényi entropies using the replica trick [41], or the ground state of a simple Kitaev chain [16, 36]. Products of these fermionic swap operators thus generate a subgroup of p-Cliffords corresponding to the symmetric group $S_{2n}\leq O(2n)$, i.e. the group of permutations. These permutations trivially preserve the operator weight of all Majorana strings.
The Cliffords corresponding to elements $S\in S_{2n}$ are a special class of fermionic Gaussian unitaries, which are defined as acting according to [42]
$U\chi_{i}\,U^{\dagger}=\sum_{j=1}^{2n}O_{ij}\chi_{j},$ (101)
where in general $O$ can be any real $2n\times 2n$ orthogonal matrix, but here must be equivalent to $S$, up to potential factors of $-1$ in some entries due to (100).
Furthermore, one can also ask which subgroup of the Gaussian Cliffords [^12] It is easy to see that non-Gaussian Cliffords can’t preserve $\hat{N}$. preserves the fermionic particle number operator
$\begin{split}\hat{N}&\coloneqq\sum_{i=1}^{n}c_{i}^{\dagger}c_{i}\\ &=\frac{1}{2}\left(n\,I+i\sum_{i=1}^{n}\chi_{2i-1}\chi_{2i}\right),\end{split}$ (102)
defined in terms of the canonical creation and annihilation operators $c_{i}\equiv(\chi_{2i-1}+i\chi_{2i})/2$ and $c_{i}^{\dagger}\equiv(\chi_{2i-1}-i\chi_{2i})/2$. Since in this case the only possible non-trivial thing the Cliffords can do is permute the terms of the sum, the symplectic action must be $S_{n}\otimes I_{2}\leq S_{2n}$. Potential sign differences such as in (100) can then be avoided due to the modes in (102) always occurring in pairs.
C.2 The Shortest Fermion Code
The shortest fermion code is a $[[6,1,3]]_{f}$ stabilizer code, meaning it utilizes 12 Majorana modes to encode one logical qubit, with the ability to correct one error and detect up to two [12]. One possible stabilizer basis for it is given by
$\begin{split}\sigma_{1}&=\chi_{1}\chi_{2}\chi_{3}\chi_{4},\\ \sigma_{2}&=\chi_{3}\chi_{4}\chi_{5}\chi_{6},\\ \sigma_{3}&=\chi_{7}\chi_{8}\chi_{9}\chi_{10},\\ \sigma_{4}&=\chi_{9}\chi_{10}\chi_{11}\chi_{12},\\ \sigma_{5}&=i\chi_{2}\chi_{4}\chi_{6}\chi_{8}\chi_{10}\chi_{12}.\\ \end{split}$ (103)
In the symplectic representation, this corresponds to the following binary basis matrix (and $v=0$):
[…]
The space of logical operators is spanned by the (unitary) Majorana strings
$\begin{split}\ell_{1}&=\chi_{1}\chi_{3}\chi_{5},\\ \ell_{2}&=\chi_{2}\chi_{4}\chi_{6}\chi_{7}\chi_{8}\chi_{9}\chi_{10}\chi_{11}\chi_{12},\end{split}$ (105)
which when multiplied together result in the total parity operator
$-\ell_{1}\ell_{2}=(-1)^{F}.$ (106)
$(-1)^{F}$ not being an element of the stabilizer implies that any two independent logical states must have opposite parity, which can be flipped by applying $\ell_{1}$ and $\ell_{2}$.
Each weight-1 error $E\sim\operatorname{\mu}(e)$ corresponding to quasiparticle poisoning can be checked against by a unique combination of stabilizers basis elements. Those combinations can be neatly read off of (104) by checking which row vectors $m$ have a non-zero entry in a given column, as for those one has $\left\langle m,e\right\rangle=m^{T}e=1$, meaning that the corresponding operators anticommute with $E$.
Since $(-1)^{F}$ is not an element of the shortest fermion code, one can use Algorithm (1) directly to get an embedding p-Clifford, without having to introduce ancillary Majorana modes. The orthogonal matrix corresponding to this Clifford is given by
[…]
Appendix D Further Proofs
D.1 Theorem 1
@@
    Using the fact that $\operatorname{\Gamma}$ is unitary, one can rewrite (14) by applying $\operatorname{\Gamma}(\cdot)\operatorname{\Gamma}^{\dagger}$ to both sides and inserting $I=\operatorname{\Gamma}^{\dagger}\operatorname{\Gamma}$:
    $\begin{split}&(\operatorname{\Gamma}\operatorname{\mu}(v)\operatorname{\Gamma}^{\dagger})(\operatorname{\Gamma}\operatorname{\mu}(v^{\prime})\operatorname{\Gamma}^{\dagger})\\ ={}&(-1)^{\left\langle v,v^{\prime}\right\rangle}(\operatorname{\Gamma}\operatorname{\mu}(v^{\prime})\operatorname{\Gamma}^{\dagger})(\operatorname{\Gamma}\operatorname{\mu}(v)\operatorname{\Gamma}^{\dagger})\end{split}$ (108)
    Evaluating both sides using (36) and dividing by $c(v)\,c(v^{\prime})\neq 0$ then leads to
    $\begin{split}&\operatorname{\mu}(S(v))\operatorname{\mu}(S(v^{\prime}))\\ ={}&(-1)^{\left\langle v,v^{\prime}\right\rangle}\operatorname{\mu}(S(v^{\prime}))\operatorname{\mu}(S(v)),\end{split}$ (109)
    On the other hand, we can also directly rewrite (14) with the substitution $v\rightarrow S(v)$:
    $\begin{split}&\operatorname{\mu}(S(v))\operatorname{\mu}(S(v^{\prime}))\\ ={}&(-1)^{\left\langle S(v),S(v^{\prime})\right\rangle}\operatorname{\mu}(S(v^{\prime}))\operatorname{\mu}(S(v)).\end{split}$ (110)
    But since both expressions have to be equivalent, we get the following requirement:
@@
    $\frac{1}{|\mathbb{F}_{2}^{2n}|}\sum_{Sv\in\mathbb{F}_{2}^{2n}}|\operatorname{tr}(\operatorname{\mu}(Sv))|^{2}=1.$ (113)
    Both representations are therefore equivalent and hence (up to an overall phase) connected by a unitary transformation $\operatorname{\Gamma}$ that induces $S$:
    $\begin{split}1&=\frac{1}{|\mathbb{F}_{2}^{2n}|}\sum_{Sv\in\mathbb{F}_{2}^{2n}}|\operatorname{tr}(\operatorname{\mu}(Sv))|^{2}\\ &=\frac{1}{|\mathbb{F}_{2}^{2n}|}\sum_{v\in\mathbb{F}_{2}^{2n}}|\operatorname{tr}(\operatorname{\Gamma}\operatorname{\mu}(v)\operatorname{\Gamma}^{\dagger})|^{2}\\ &=\frac{1}{|\mathbb{F}_{2}^{2n}|}\sum_{v\in\mathbb{F}_{2}^{2n}}|\operatorname{tr}(\operatorname{\mu}(v)\operatorname{\Gamma}^{\dagger}\operatorname{\Gamma})|^{2}\\ &=\frac{1}{|\mathbb{F}_{2}^{2n}|}\sum_{v\in\mathbb{F}_{2}^{2n}}|\operatorname{tr}(\operatorname{\mu}(v))|^{2}.\end{split}$ (114)
@@
    $\begin{split}&(\operatorname{\mu}(a)\operatorname{\Gamma}(S))\operatorname{\mu}(v)(\operatorname{\mu}(a)\operatorname{\Gamma}(S))^{\dagger}\\ ={}&\operatorname{\mu}(a)\operatorname{\Gamma}(S)\operatorname{\mu}(v)\operatorname{\Gamma}^{\dagger}(S)\operatorname{\mu}(a)\\ ={}&c(v)\cdot\operatorname{\mu}(a)\operatorname{\mu}(Sv)\operatorname{\mu}(a)\\ ={}&(-1)^{\left\langle Sv,a\right\rangle}c(v)\cdot\operatorname{\mu}(Sv)\\ \equiv{}&c^{\prime}(v)\cdot\operatorname{\mu}(Sv),\end{split}$ (115)
@@
    $\left(\frac{c(v)}{c^{\prime}(v)}\right)\cdot\left(\frac{c(v^{\prime})}{c^{\prime}(v^{\prime})}\right)=\frac{c(v+v^{\prime})}{c^{\prime}(v+v^{\prime})}$ (116)
    for all $v,v^{\prime}\in\mathbb{F}_{2}^{2n}$. This together with $c(v)=\pm 1$ (and the symplectic form being non-degenerate) implies that – without loss of generality – $c(v)/c^{\prime}(v)$ can be written as
    $\frac{c(v)}{c^{\prime}(v)}\equiv(-1)^{\left\langle v,a^{\prime}\right\rangle}=(-1)^{\left\langle Sv,a\right\rangle},$ (117)
    for some $a^{\prime}\equiv S^{-1}a\in\mathbb{F}_{2}^{2n}$ due to (38). This coincides with the form of the phase induced by an additional Majorana string $\operatorname{\mu}(a)$, as shown in (LABEL:eq:cliff_majo_action).
@@
    $\begin{split}&\operatorname{\Gamma}(S)\operatorname{\Gamma}(T)\operatorname{\mu}(v)\operatorname{\Gamma}^{\dagger}(T)\operatorname{\Gamma}^{\dagger}(S)\\ \propto{}&\operatorname{\Gamma}(S)\operatorname{\mu}(Tv)\operatorname{\Gamma}^{\dagger}(S)\\ \propto{}&\operatorname{\mu}(STv)\\ \propto{}&\operatorname{\Gamma}(ST)\operatorname{\mu}(v)\operatorname{\Gamma}^{\dagger}(ST)\quad\forall v\in\mathbb{F}_{2}^{2n},\end{split}$ (118)
@@ D.2 Theorem 5
Next we first show that $\lvert\operatorname{tr}(\Pi_{+}U)\rvert^{2}$ with $U=\operatorname{\mu}(a)\operatorname{\Gamma}(S)\in\mathcal{G}$ is either zero or equal to $(f_{+}(S)+c_{+}(S))/2$. For that we evaluate the expression as a sum over the Bell basis
$\lvert\Phi_{n}(v)\rangle\equiv(\operatorname{\mu}(v)\otimes I)\lvert\Phi_{n}^{+}\rangle\in\operatorname{\mathcal{H}}\otimes\operatorname{\mathcal{H}},$ (120)
where $\lvert\Phi_{n}^{+}\rangle\equiv\lvert\Phi^{+}\rangle^{\otimes n}$ denotes $n$ copies of canonical Bell state. That these states indeed form a complete basis of $\operatorname{\mathcal{H}}\otimes\operatorname{\mathcal{H}}$ follows from:
@@
$\begin{split}\langle\Phi_{n}(v)|\Phi_{n}(v^{\prime})\rangle&=\langle\Phi_{n}^{+}|\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})\otimes I|\Phi_{n}^{+}\rangle\\ &=\frac{1}{2^{n}}\operatorname{tr}(\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime}))\\ &=\delta_{v,v^{\prime}},\end{split}$ (121)
(using (34) and a well-known property of the canonical Bell state) and the fact that $\dim(\operatorname{\mathcal{H}}\otimes\operatorname{\mathcal{H}})=2^{2n}=|\mathbb{F}_{2}^{2n}|$. With that we can evaluate the aforementioned trace:
$\begin{split}&\lvert\operatorname{tr}(\Pi_{+}U)\rvert^{2}\\ ={}&\operatorname{tr}(\Pi_{+}U\otimes\Pi_{+}U^{\dagger})\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}\langle\Phi_{n}(v)\rvert\Pi_{+}U\otimes\Pi_{+}U^{\dagger}\lvert\Phi_{n}(v)\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\Pi_{+}U\operatorname{\mu}(v)\otimes\Pi_{+}U^{\dagger}\lvert\Phi_{n}^{+}\rangle\\ \stackrel{{\scriptstyle\text{(a)}}}{{=}}{}&\sum_{v\in\mathbb{F}_{2}^{2n}}\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\Pi_{+}U\operatorname{\mu}(v)U^{\dagger}\Pi_{+}\otimes I\lvert\Phi_{n}^{+}\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}c(v)\,\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\Pi_{+}\operatorname{\mu}(Sv)\Pi_{+}\otimes I\lvert\Phi_{n}^{+}\rangle\\ \stackrel{{\scriptstyle\text{(b)}}}{{=}}{}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}c(v)\,\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\operatorname{\mu}(Sv)\Pi_{+}\otimes I\lvert\Phi_{n}^{+}\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{c(v)}{2^{n}}\,\operatorname{tr}[\operatorname{\mu}(v)\operatorname{\mu}(Sv)\Pi_{+}]\\ \stackrel{{\scriptstyle\text{(c)}}}{{=}}{}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{c(v)}{2^{n+1}}\,\operatorname{tr}[\operatorname{\mu}(v)\operatorname{\mu}(Sv)+\operatorname{\mu}(v)\operatorname{\mu}(Sv)\operatorname{\mu}(j)]\\ \stackrel{{\scriptstyle\text{(d)}}}{{=}}{}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{c(v)}{2}\,(\delta_{Sv,v}+\zeta(j,v)\,\delta_{Sv,v^{c}}).\end{split}$ (122)
The steps labeled in the derivation are achieved in the following ways:
  1. (a)
    The equivalence is straightforward to prove by expressing both inner products using circuit notation (also called Penrose graphical notation), akin to how it is done for quantum teleportation.
  2. (b)
    $\Pi_{+}\operatorname{\mu}(Sv)\Pi_{+}$ either vanishes if $Sv$ has odd parity, or is equal to $\operatorname{\mu}(Sv)\Pi_{+}$ if $Sv$ has even parity. And since $S$ preserves parity, this means that the sum over $v\in\mathbb{F}_{2}^{2n}$ reduces to the sum over $v\in\mathbb{F}_{2}^{2n,+}$.
  3. (c)
    We insert the definition (85) of the projector $\Pi_{+}$ with the identification $(-1)^{F}\sim\operatorname{\mu}(j).$
  4. (d)
    We fuse $\operatorname{\mu}(v)$ and $\operatorname{\mu}(j)$ using trace cyclicity and (7), and then apply (34) to both terms.
We can now analyze the structure of $c(v)$ for the resulting cases of $Sv=v$ and $Sv=v^{c}$:
@@
  • In the case of $Sv=v$, it follows directly from (44) that the coefficients of any two (even-parity) fixed points $v,v^{\prime}$ satisfy $c(v)\cdot c(v^{\prime})=c(v+v^{\prime})$. This together with (45) and (LABEL:eq:cliff_majo_action) implies that we can identify $c(v)\equiv(-1)^{\left\langle Sv,a\right\rangle}=(-1)^{a^{T}v}$, up to an overall sign that might arise to make the relevant terms in (LABEL:eq:trace_squared_parity_1) positive, but which can be ignored without loss of generality.
  • In the case of $Sv=v^{c}$, (44) can be rewritten as
    $c(v)\cdot c(v^{\prime})=\frac{\zeta(v,v^{\prime})}{\zeta(v^{c},(v^{\prime})^{c})}\cdot c(v+v^{\prime}),$ (123)
@@
    for any two (even-parity) complemented points $v,v^{\prime}$. But since the sum of two such points is always a fixed point, we can identify $c(v+v^{\prime})=(-1)^{a^{T}(v+v^{\prime})}$ as before. Additionally, by evaluating $\operatorname{\mu}(j)\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})\operatorname{\mu}(j)$ in two different orders one can arrive at
    $\zeta(v,v^{\prime})=\zeta(j,v)\cdot\zeta(j,v^{\prime})\cdot\zeta(v^{c},(v^{\prime})^{c}).$ (124)
@@
    Plugging this all into (123) and using $\zeta(j,v)=\pm 1$ (since $\left\langle j,v\right\rangle=0$), results in
    $\left(c(v)\,\zeta(j,v)\right)\cdot\left(c(v^{\prime})\,\zeta(j,v^{\prime})\right)=(-1)^{a^{T}(v+v^{\prime})},$ (125)
@@
Applying these results to (LABEL:eq:trace_squared_parity_1) then leaves us with